The length of a standard jewel case is 5 cm more than its width. The area of the rectangular top of the case is 266 cm squared.

What is the length and width of the jewel case?

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

width=x
length=x+5
area is x(x+5)=266
so x^2+5x-266=0
(x-14)(x-19)=0
x=14 cm width
x+5=19 cm length

Answer by Alan3354(69443)

The length of a standard jewel case is 5 cm more than its width. The area of the rectangular top of the case is 266 cm squared.

What is the length and width of the jewel case?
--------
Find a pair of factors of 266 that differ by 5.
----
Or, if you prefer to spend more time on it:
L = W + 5
L*W = 266
W*(W+5) = 266
W^2 + 5W = 266
W^2 - 5W - 266 = 0
Now, find a pair of factors of 266 that differ by 5.